### v1

# Word Order in English-Japanese Simultaneous Interpretation: Analyses and Evaluation using Chunk-wise Monotonic Translation

Abstract
This paper analyzes the features of monotonic translations, which follow the word order of the source language, in simultaneous interpreting (SI). The word order differences are one of the biggest challenges in SI, especially for language pairs with significant structural differences like English and Japanese. We analyzed the characteristics of monotonic translations using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset and found some grammatical structures that make monotonic translation difficult in English-Japanese SI. We further investigated the features of monotonic translations through evaluating the output from the existing speech translation (ST) and simultaneous speech translation (simulST) models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing test sets. The results suggest that the existing SI-based test set underestimates the model performance. We also found that the monotonic-translation-based dataset would better evaluate simulST models, while using an offline-based test set for evaluating simulST models underestimates the model performance.
Word Order in English-Japanese Simultaneous Interpretation:
Analyses and Evaluation using Chunk-wise Monotonic Translation
Kosuke Doi1, Yuka Ko1, Mana Makinae1, Katsuhito Sudoh1, 2, Satoshi Nakamura1, 3 1Nara Institute of Science and Technology, 2Nara Women’s University, 3The Chinese University of Hong Kong, Shenzhen {doi.kosuke.de8, sudoh, s-nakamura}@is.naist.jp
1 Introduction
Simultaneous Interpreting (SI) is the task of translating speech from a source language into a target language in real time. SI is a cognitively demanding task, and human simultaneous interpreters are known to employ strategies such as segmentation, summarization, and generalization He et al. (2016). Maintaining the word order in a source language is also an important strategy, especially for language pairs whose word order is different (e.g., English and Japanese), to shorten delay and reduce cognitive load. Because of these features, SI sentences are different from offline translation sentences, but most of the studies in automatic SI Oda et al. (2014); Ma et al. (2019); Liu et al. (2020); Papi et al. (2023) have been used offline translation corpora (e.g., MuST-C; Di Gangi et al., 2019) for both training and evaluation of the models due to the limited amount of simultaneous interpretation corpora (SICs).
For the English-Japanese language pair, several SICs have been constructed Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021). Based on NAIST Simultaneous Interpretation Corpus (NAIST-SIC; Doi et al., 2021), Zhao et al. (2024)111The dataset was released in 2023 (see the version 3 of the paper). created an automatically-aligned parallel SI dataset, named NAIST-SIC-Aligned. Since the sentences in the dataset are aligned at the sentence level, they can be used for model training. Actually, Ko et al. (2023) and Zhao et al. (2024) trained SI models using SI data from NAIST-SIC-Aligned. The model performance was evaluated through automatic evaluation metrics such as BLEU Papineni et al. (2002) using a small test set curated based on SI sentences generated by professional human simultaneous interpreters.
The scores reported in Ko et al. (2023) and Zhao et al. (2024) were relatively low but the test set used in the two studies might have underestimated the model performance. Since human simultaneous interpreters use strategies such as summarization and generalization, phrases that do not affect the main idea are not necessarily translated into the target language. If an SI model generates translation for the phrases that a human interpreter did not, the output sentence might not be evaluated properly, even when it is correct translation.
| Source | (1) The US Secret Service, / (2) two months ago, / (3) froze the Swiss bank account / (4) of Mr. Sam Jain right here, / (5) and that bank account / (6) had 14.9 million US dollars in it / (7) when it was frozen. |
| Offline | {CJK} UTF8min (1) 米国のシークレットサービスは / (2) 2ヶ月前に / (4) サム・ジェイン氏の / (3) スイス銀行口座を凍結しました / (5) その口座には / (6) 米ドルで1490万ドルありました |
| [The US Secret Service / two months ago / Mr. Sam Jain’s / froze the Swiss bank account / that bank account / had 14.9 million US dollars] | |
| SI | {CJK} UTF8min (1) アメリカのシークレッドサービスが、/ (3) スイスの銀行の口座を凍結しました。 / (4) サムジェインのものです。 / (5) この銀行口座の中には、 / (6) 一千四百九十万ドルが入っていました。 |
| [The US Secret Service / froze the Swiss bank account / it is Sam Jain’s one / in this bank account / had 14.9 million dollars] | |
| CMT | {CJK} UTF8min (1) アメリカ合衆国シークレットサービスは、 / (2) 2ヶ月前に、 / (3) スイスの銀行口座を凍結しました、 / (4) ここにいるサム・ジェイン氏の口座です、 / (5) そしてその銀行口座には / (6) 490万米ドルが入っていました、 / (7) 凍結された時。 |
| [The US Secret Service / two months ago / froze the Swiss bank account / the account of Mr. Sam Jain right here / and that bank account / had 14.9 million US dollars in it / when it was frozen] |
Fukuda et al. (2024) pointed out the difficulty for SI models to learn which phrases in source speech are less important and called for constructing SI models that only employ the strategy of maintaining the word order in a source language. As the first step, they created NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset222https://dsc-nlp.naist.jp/data/NAIST-SIC/Aligned-Chunk_Mono-EJ/. The source sentences in the test set used in Ko et al. (2023) were automatically segmented into chunks, and each chunk was translated in a way that did not include the content of the subsequent chunks. Unlike in SI sentences by human interpreters, where not all information in the source sentences is translated, chunk-wise monotonic translation (CMT) was developed so that all information is translated (Table 1)333Precisely, omission to maintain the fluency of the sentence was allowed. See Section 3.1 for the details about the dataset.. Fukuda et al. (2024) have investigated the quality of CMT through human evaluation, but have not analyzed the characteristics of it. In addition, they have not conducted any evaluation experiments in which model outputs are evaluated on the CMT Evaluation dataset.
In this paper, we analyze CMT sentences qualitatively and quantitatively. In the process of developing CMT, it was allowed to repeat, defer, and omit phrases in the source sentences to keep the translation fluent. We suppose that there are factors (e.g., syntactic structures) that prevent monotonic translation if phrases were repeated, deferred, or omitted in CMT sentences since the sentences were translated without time constraints. In addition, we evaluate the output from (1) a speech translation (ST) model trained on offline data, (2) a simultaneous speech translation (simulST) model trained on offline data, and (3) a simulST model trained on both offline and SI data. The contributions of this paper are as follows:
We analyze CMT sentences and show that they tend to be longer than offline translation primarily because of repetition.
We investigate what makes the phrases in source sentences to be repeated, deferred, and omitted, and show that most cases occur because of particular grammatical structures. When a phrase in a chunk is a dependent of a phrase in the preceding chunk, the head phrase is typically repeated or deferred.
We evaluate the output from three different models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset: (1) an ST model trained on offline data, (2) a simulST model trained on offline data, and (3) a simulST model trained on both offline and SI data. We show that the existing SI-based test set Ko et al. (2023); Zhao et al. (2024) underestimates the model performance. The results also show that the CMT evaluation dataset would better evaluate simulST models while using an offline-based test set for evaluating simulST models underestimates the model performance.
2 Related Work
2.1 Simultaneous Interpretation Corpora
SICs are valuable resources both for developing automatic SI models and for analyzing the characteristics of SI. For the English-Japanese language pair, several SICs are publicly available Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021) although the amount of such corpora is very limited compared to offline translation corpora.
Using these corpora, SI sentences have been analyzed from various perspectives, such as strategies and interpreting patterns used by interpreters, latency, translation quality, and word order Tohyama and Matsubara (2006); Ono et al. (2008); Cai et al. (2018, 2020); Doi et al. (2021). Furthermore, SI models have also been developed using SICs Ryu et al. (2004); Shimizu et al. (2013); Ko et al. (2023).
2.2 Word Order in Simultaneous Interpreting
When dealing with language pairs whose sentence structures are different, including English and Japanese (SVO and head-initial vs. SOV and head-final), reducing the word order difference between the source and the target language is important to minimize the delay.
Murata et al. (2010) segmented source sentences into semantically meaningful units with a maximum length of 4.3 seconds, and translated those units from the viewpoint of SI. He et al. (2015) designed syntactic transformation rules for Japanese-English simultaneous machine translation. By applying the rules to target language sentences (i.e., English), they generated more monotonic translations, while preserving the meaning of source sentences and maintaining the grammaticality of the target language. In English-Japanese SI, Futamata et al. (2020) reordered Japanese sentences to make the word order closer to the original English sentences. They further applied style transfer to make the sentence fluent and obtained sentences close to SI sentences by human interpreters. Han et al. (2021) proposed an algorithm to reorder and refine the target sentences so that the target sentences were aligned largely monotonically. They trained SI models for four language pairs, including English-Japanese. Nakabayashi and Kato (2021) segmented sentences into chunks and created bilingual pairs of the chunks with explicit annotations of context information. The SI model trained on the data translated the source sentences while referencing preceding chunks but connecting chunks naturally remained a challenge. Higashiyama et al. (2023) have constructed a large-scale English Japanese SIC with information of chunk boundaries in source and target sentences and phrases that can be omitted in target sentences. NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset Fukuda et al. (2024), which is similar to Higashiyama et al.’s (2023), is relatively small and intended for the evaluation purpose.
The word order differences among different translation modes have also been investigated. Okamura and Yamada (2023) quantitatively compared the degree to which the word order of the source sentences was maintained and found that SI sentences retained the order better than consecutive interpreting and offline translation sentences. Cai et al. (2020) found syntactic and non-syntactic factors that affect interpreters’ word order decisions through the statistical analyses of an SIC. In this paper, we analyze what makes monotonic translation difficult. While Cai et al. (2020) analyzed actual SI data generated by human simultaneous interpreters, we use the CMT dataset, which is created without time constraints, and in which all information in the source sentences is translated into the target sentences.
3 Chunk-wise Monotonic Translation
In SI between language pairs with different sentence structures, interpreters are known to segment source sentences into chunks and translate them in the order of the chunks444Reordering may occur within a chunk. Okamura and Yamada (2023). This section describes the details of NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset, which is used in our analyses.
3.1 Data
NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset consists of 511 pairs of source sentences and their corresponding chunk-wise monotonic translation (CMT) sentences with information of chunk boundaries in source and target sentences. Source (i.e., English) sentences, which were used as the test set in Ko et al. (2023), were segmented following the five rules that reflected interpreter’s strategies555See the original paper or the code for chunking: https://github.com/ahclab/si_chunker based on the syntactic analysis results from spaCy. The source sentences come from 8 TED talks.
| Data | Sum | Per Sent. |
| # sentence pairs | 511 | – |
| # chunks | 1,677 | 3.28 |
| # source words | 8,104 | 15.86 |
| # target words | 13,981 | 27.36 |
Translators were provided with source sentences with chunk boundaries and were asked to (1) translate them in the order of the chunks while (2) connecting the chunks naturally and (3) not including the content of the subsequent chunks. It was allowed to (1) repeat, (2) defer, and (3) omit phrases in the source sentences to keep the translation fluent, but was instructed to avoid using the operations with larger number as much as possible. Data statistics are shown in Table 2.
4 Data Analysis
Fukuda et al. (2024) have examined the quality of CMT sentences through human evaluation but have not analyzed the characteristics of them. We suppose that there are factors that prevent monotonic translation if a phrase in the source sentences is repeated, deferred, and omitted since the CMT sentences were generated without time constraints. We qualitatively and quantitatively analyze the CMT sentences with these operations and reveal the factors.
To better understand the characteristics of CMT sentences, we also compare them with SI sentences from NAIST-SIC and NAIST-SIC-Aligned, and offline translation sentences from the subtitles of TED talks. Since the SI sentences in NAIST-SIC were not aligned at the sentence level, we performed alignment manually. Some source sentences did not match across the datasets, and we excluded those 10 sentences from the analyses. In addition, 25 sentences were not translated in NAIST-SIC666For example, due to time constraints, interpreters may not be able to translate a whole sentence. See Doi et al. (2021)., which were also excluded from the analyses. As a result, 476 sentences were used for our analyses.
| Type | Tag | Meaning |
| Span | repeat | Phrases that are repeated |
| zero-repeat | Target phrases that are repeated; No corresponding source phrases (e.g., zero that-clause) | |
| defer | Phrases that are not translated within the current chunk but in a subsequent chunk | |
| omit | Source phrases that are not translated in the target sentences | |
| ahead | Target phrases translated using the subsequent chunks | |
| add | Target phrases that have no corresponding source phrases | |
| error | Phrases with translation errors | |
| sep | Boundaries of source and target sentences | |
| Relation | rel-repeat | Connect source and target phrases with repeat tags |
| repeat_d# | Connect target phrases with repeat tags | |
| defer_d# | Connect source and target phrases with defer tags | |
| ahead_d# | Connect source and target phrases with ahead tags | |
| rel-err | Connect source and target phrases with error tags |
4.1 Annotations
To analyze the characteristics of CMT sentences, we annotated tags to source and CMT sentences. The list of tags is shown in Table 3. Prior to the annotation, we tokenized the English sentences using spaCy777https://spacy.io/ and the Japanese sentences using MeCab Kudo et al. (2004) with unidic. Then, the source and CMT sentences were concatenated with special token [SEP] and annotated using an open-source data labeling tool, doccano888https://github.com/doccano/doccano. Annotation examples are shown in Figure 1.
The first and second authors collaboratively annotated the first 50 examples while discussing. As sufficient agreement was expected, the remainder of the data were annotated by the first author.
4.2 Results of Analyses
4.2.1 Comparison among Different Translation Modes
We compared the length of sentences in the four datasets. Fukuda et al. (2024) have also conducted similar comparisons based on the number of characters. However, there were variations in spelling and differences in transcription systems (e.g., numbers), and we made comparisons based on the number of words segmented by MeCab.
Table 4 show that CMT sentences were the longest, followed by offline, NAIST-SIC, and NAIST-SIC-Aligned, which matched the results reported in Fukuda et al. (2024). Long translation sentences can pose some problems. As discussed in Fukuda et al. (2024), it may increase the cognitive load on the listeners/readers of the translations. In addition, longer output may cause a delay even though CMT aims to reduce it.
| Dataset | Sum | Per Sent. |
| CMT | 13,508 | 28.38 |
| NAIST-SIC | 8,914 | 18.73 |
| NAIST-SIC-Aligned | 8,072 | 16.96 |
| Offline | 9,907 | 20.81 |
4.2.2 Factors that Make CMT Sentences Longer
To reveal what made the CMT sentences longer, we first analyzed them qualitatively. Our analyses suggest that (1) CMT sentences contain many formulaic expressions for the end of sentences as sentences are segmented into small chunks, and (2) words that are often omitted in Japanese (e.g., pronouns) are explicitly translated since translators were instructed to avoid omitting phrases in the source language, as in the following examples.
(En) It’s when we warmed it up, / and we turned on the lights / and looked inside the box, / we saw that the piece metal / was still there in one piece.
{CJK}UTF8min (CMT) 私たちがそれを暖めるときです、 / 電気をつけて、 / 箱の中を見たときです、 / 私たちはその金属片を見たんです、 / それはまだ一つの塊としてそこにありました。
(offline) 物体を暖め明かりをつけ 箱の中を見たところ金属片はまだそこに存在していました
(En) He only has use of his eyes.
{CJK}UTF8min (CMT) 彼は目だけを使えます。
(offline) 目だけしか動かせません
In addition to the above characteristics, there were many repetitions observed in particularly long sentences. To verify this, we further analyzed particularly long and short sentences, chosen based on the length ratio of CMT sentences to offline ones. The long and short sentences were defined as those with a length ratio greater/smaller than the average 0.5 standard deviations (avg.=1.39, SD=0.43). We subjectively judged if those sentences contained many repetitions. We also identified the sentences whose offline translations were short.
Table 5 shows that long sentences contained more repetitions than short ones. The offline translation sentences were short, probably because they were originally subtitles, for which limited space was allowed. We also checked it quantitatively using the number of repeat tags assigned to the sentences, and found that the frequency of repetition tags was higher in long sentences (Table 5).
| Type | N | Repeat | Repeat | Short |
| (subjective) | (tag) | offline | ||
| Long | 131 | 54 (41.22%) | 3.35 | 53 |
| Short | 171 | 22 (12.87%) | 0.81 | – |
4.2.3 Omission in SI Sentences
To find techniques to shorten translations, we analyzed SI sentences in NAIST-SIC. Based on the length ratio of SI sentences to offline ones, we defined SI sentences that potentially involved reasonable omissions (omission; the ratio ), and SI sentences that may not fully convey the meaning of the source sentences (undertranslation; the ratio ), following the criteria in Higashiyama et al. (2023).
We expected to find some trends (e.g., part-of-speech) in phrases that were omitted, but we was not able to. In addition, we found a certain number of unacceptable translations in both categories (43.12% and 60.00% for omission and undertranslation, respectively). The results suggest that human simultaneous interpreters judge the importance of the phrases based on context and decide whether to translate them, which are sometimes correct and sometimes not.
4.2.4 Factors that Make Monotonic Translation Difficult
With the help of tags annotated to source and CMT sentences, we analyzed factors that make monotonic translation difficult. Table 6 shows the number of source phrases that were repeated, deferred, and omitted. The values are based on the number of rel-repeat, defer_d#, and omit tags, respectively. We counted the number of relation tags for repeat and defer because the span tags for the two operations were assigned to both source and CMT sentences. The results show that the translators used repeat most frequently, followed by defer and omit, as they were instructed (see Section 3.1).
| Operation | N |
| repeat | 301 |
| defer | 173 |
| omit | 36 |
For these phrases, we explored what makes monotonic translations difficult. Our analyses revealed that most cases of repeat and defer occurred because of particular grammatical structures. Table 7 lists major structures along with their frequencies in the data and examples. In these structures, a phrase in a chunk is typically a dependent of a phrase in the preceding chunk. In the example of a post-modifier (Table 7), the relative pronoun clause is a dependent of the noun phrase a device, which is in a preceding chunk. When phrases with a dependency relation exist across multiple chunks, CMT would be difficult because Japanese is a strongly head-final language. The examples in Table 7 show how human translators address these structures by repeating or deferring some phrases in the subsequent chunks.
Prepositions, post-modifiers, and dependent clauses have also been identified as syntactic factors that affect interpreters’ word order decisions in Cai et al. (2020). Human interpreters find those structures challenging for SI and adopt a strategy to maintain the word order of the source language.
| Structures | # repeat | # defer | Examples |
| Noun with a post-modifier | 88 | 12 | And now we’ve created a device / that has absolutely no limitations. |
| {CJK} UTF8min さて、私たちはデバイスを作り出しました、 / 全く制限のないものです。 [And we’ve created a device / one that has absolutely no limitations] | |||
| Head followed by multiple dependents | 35 | 6 | … / allows for deep squats, / crawls and high agility movements. |
| {CJK} UTF8min … / 深いスクワットを可能にし、 / クロールや高い敏捷性の動きを可能にします。 [… / allows for deep squats / allows for crawls and high agility movements] | |||
| Dependent conjunction | 26 | 6 | … / when he’s covered / in four feet of snow. |
| {CJK} UTF8min … / 彼が / 四フィートの雪に覆われてしまっていた時には。 [he / when was covered in for feet of snow] | |||
| Chunk boundary before a clause | 15 | 13 | … / you know that this isn’t / how it normally goes. |
| {CJK} UTF8min / あなたは分かるはずです、これは / 通常の進行ではないということが。 [you know that this / isn’t how it normally goes] | |||
| Chunk boundary before a preposition | 10 | 7 | … / providing totalitarian governments with tools / to do this / against their own citizens. |
| {CJK} UTF8min … / 全体主義政府にツールを提供しているということです、 / これを行うための、 / 自国の市民に対してこれを行うためのツールを。 [providing totalitarian governments with tools / to do this / tools to do this against their own citizens] |
In addition, we observed that inappropriate segmentation was addressed through repeating and deferring the phrases. Most of the inappropriate segmentation was found in phrasal verbs, -ing forms of verbs, and to-infinitives.
In SI data, we also found human interpreters repeat phrases to maintain the word order of the source language. For example, in the example in Table 1, a noun modified by a preposition phrase is repeated.
(En) … / froze the Swiss bank account / of Mr. Sam Jain right here, / …
{CJK}UTF8min (SI) … / スイスの銀行の口座を凍結しました。 / サムジェインのものです。 / …
[… / froze the Swiss bank account / it is Sam Jain’s one / …]
In addition, Okamura and Yamada (2023) reported that the order of the chunks is swapped about once on average in an SI sentence. These things suggest that human interpreters address the word order differences that make monotonic translation difficult through repeating and deferring some phrases.
5 Evaluation Using CMT Data
To investigate the impact of using CMT sentences for evaluating the translation quality, we evaluated the output from existing ST and simulST models using NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as existing test sets.
5.1 Data
We used the following datasets as the reference for the automatic evaluation metrics:
n-cmt: CMT sentences from NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset
si_hum: SI sentences from NAIST-SIC, which were manually aligned to the source speech
si_auto: SI sentences from NAIST-SIC-Aligned, which were aligned automatically
offline: offline translation sentences from the subtitles of TED talks.
5.2 Speech Translation Models
We used three existing ST and simulST models:
ST_offline: an ST model trained on offline data Fukuda et al. (2023)
simulST_offline: a simulST model trained on offline data Ko et al. (2023)
simulST_si_offline: a simulST model trained on both offline and SI data Ko et al. (2023)
All the models were built by connecting two pre-trained models, HuBERT-Large Hsu et al. (2021) for their speech encoder and the decoder of mBART50 Tang et al. (2020) for their text decoder. The encoder and decoder were connected via Inter-connection Nishikawa and Nakamura (2023) and a length adapter Tsiamas et al. (2022). Both SimulST models used bilingual prefix pairs extracted using Bilingual Prefix Alignment Kano et al. (2022) for the model training and employed a decoding policy called local agreement Liu et al. (2020). For ST_offline, we used the model with checkpoint averaging (Inter-connection + Ckpt Ave. in Fukuda et al. (2023)). For simulST_offline and simulST_si_offline, we used the models which satisfy the task requirement of the simultaneous track in IWSLT 2023 Evaluation Campaign999https://iwslt.org/2023/simultaneous, latency measured by Average Lagging Ma et al. (2019) seconds (Offline FT and Mixed FT + Style in Ko et al. (2023), respectively).
5.3 Metrics
We evaluated the translation quality of the output from the ST and simulST models (see Section 5.2) using BLEU101010BLEU was calculated using sacreBLEU. Post (2018) https://github.com/mjpost/sacrebleu, BLEURT111111https://github.com/google-research/bleurt Sellam et al. (2020), COMET121212https://github.com/Unbabel/COMET Rei et al. (2020), and BERTScore131313https://github.com/Tiiiger/bert_score Zhang et al. (2020). BERTScore was calculated using bert-base-multilingual-cased. We used the four datasets described in Section 5.1 as the reference.
5.4 Evaluation Results
| Model | BLEU | BLEURT | COMET | |||||||||
| n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | |
| ST_offline | 14.487 | 8.856 | 8.637 | 17.775 | 0.553 | 0.447 | 0.414 | 0.538 | 0.838 | 0.797 | 0.781 | 0.833 |
| simulST_offline | 15.406 | 8.446 | 7.773 | 17.907 | 0.556 | 0.442 | 0.406 | 0.531 | 0.826 | 0.780 | 0.763 | 0.821 |
| simulST_si_offline | 15.982 | 12.031 | 11.020 | 13.191 | 0.567 | 0.493 | 0.460 | 0.519 | 0.807 | 0.774 | 0.761 | 0.789 |
| Model | BERTScore (Pre.) | BERTScore (Rec.) | BERTScore (F1) | |||||||||
| n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | |
| ST_offline | 0.801 | 0.735 | 0.722 | 0.789 | 0.769 | 0.739 | 0.735 | 0.788 | 0.784 | 0.737 | 0.728 | 0.788 |
| simulST_offline | 0.799 | 0.730 | 0.717 | 0.783 | 0.770 | 0.738 | 0.734 | 0.786 | 0.783 | 0.734 | 0.725 | 0.784 |
| simulST_si_offline | 0.817 | 0.764 | 0.746 | 0.759 | 0.784 | 0.766 | 0.760 | 0.757 | 0.800 | 0.764 | 0.752 | 0.757 |
Table 8 shows the results of quality evaluation metrics across ST and simulST models. When the translation quality was evaluated using BLEU with n-cmt as the reference, simulST_si_offline achieved the highest score. On the SI-based test sets (i.e., si_hum and si_auto), simulST_si_offline also had the highest score. On the offline-based test set, in contrast, models trained on only offline data achieved much higher score than simulST_si_offline. The same tendencies were observed in BLEURT and BERTScore. These results suggest that the model trained on both SI and offline data generated more SI-like translations, and that such models should be evaluated using a reference closer to SI sentences. In addition, using an offline-based test set might underestimate the performance of the model trained on both SI and offline data.
Comparing n-cmt, si_hum, and si_auto, the scores were highest for n-cmt, followed by si_hum and si_auto on all metrics and models. Because si_hum is based on SI sentences generated by human simultaneous interpreters, some content in a source speech might be omitted or inadequately translated (under-translation). SI sentences in si_auto, which were automatically created based on human SI sentences, might contain less source speech content than those in si_hum due to the alignment and filtering techniques applied (see Zhao et al., 2024). In fact, BERTScore precision was higher than recall on n-cmt, in which there are almost no omissions, while the recall was higher than the precision on si_auto and the precision and the recall were almost equal on si_hum. These results suggest that the existing SI-based test sets Ko et al. (2023); Zhao et al. (2024) underestimate the model performance.
However, the results on COMET were different from those on other metrics (Table 8). On all four test sets, ST_offline achieved the highest score, followed by simulST_offline and simulST_si_offline. One of the possible reasons is that COMET uses source sentences to calculate scores.
To examine the impact of source sentences, we also calculated the reference-free COMET-QE using wmt22-cometkiwi-da and got similar results (0.813, 0.798, 0.766 for ST_offline, simulST_offline, simulST_si_offline, respectively). We further calculated COMET-QE for n-cmt and offline, regarding them as oracle data, and found that n-cmt had a higher score than offline (n-cmt: 0.832, offline: 0.812). Because some translation sentences in offline include under-translation, these results suggest that the COMET score tends to become higher when more content in the source sentences is covered in the target sentences. This feature does not fit the nature of SI, where human interpreters use sophisticated strategies (see He et al., 2016; Cai et al., 2020, for example). We need to carefully interpret the COMET score when we use it for evaluating simulST models.
6 Conclusion
This paper focuses on monotonic translations in English-Japanese SI. Our analyses revealed some syntactic factors that make monotonic translations difficult and that human interpreters/translators address these challenges by repeating or deferring some phrases in the source language in the subsequent chunks. The grammatical structures that might cause delay would be useful information for developing segmentation or decoding policies of simultaneous machine translation systems.
We also evaluated the output from the existing ST and simulST models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing SI-based and offline-based test sets. The results suggest that monotonic translation data would better evaluate simulST models trained using SI data, and that the existing test sets underestimate the performance of such models. A future direction would be using monotonic translation data for developing simulST models.
Acknowledgments
A part of this work was supported by JSPS KAKENHI Grant Numbers JP21H05054 and by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2137.

### v2

@@
 This paper analyzes the features of monotonic translations, which follow the word order of the source language, in simultaneous interpreting (SI).
-The word order differences are one of the biggest challenges in SI, especially for language pairs with significant structural differences like English and Japanese.
-We analyzed the characteristics of monotonic translations using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset and found some grammatical structures that make monotonic translation difficult in English-Japanese SI.
-We further investigated the features of monotonic translations through evaluating the output from the existing speech translation (ST) and simultaneous speech translation (simulST) models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing test sets.
-The results suggest that the existing SI-based test set underestimates the model performance.
-We also found that the monotonic-translation-based dataset would better evaluate simulST models, while using an offline-based test set for evaluating simulST models underestimates the model performance.
+Word order differences are one of the biggest challenges in SI, especially for language pairs with significant structural differences like English and Japanese.
+We analyzed the characteristics of chunk-wise monotonic translation (CMT) sentences using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset and identified some grammatical structures that make monotonic translation difficult in English-Japanese SI.
+We further investigated the features of CMT sentences by evaluating the output from the existing speech translation (ST) and simultaneous speech translation (simulST) models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing test sets.
+The results indicate the possibility that the existing SI-based test set underestimates the model performance.
+The results also suggest that using CMT sentences as references gives higher scores to simulST models than ST models, and that using an offline-based test set to evaluate the simulST models underestimates the model performance.
 Word Order in English-Japanese Simultaneous Interpretation:
@@
 Simultaneous interpreting (SI) is the task of translating speech from a source language into a target language in real time.
-SI is a cognitively demanding task, and human simultaneous interpreters are known to employ strategies such as segmentation, summarization, and generalization He et al.
+SI is cognitively demanding, and human simultaneous interpreters employ such strategies as segmentation, summarization, and generalization He et al.
 (2016).
-Maintaining the word order in a source language is also an important strategy, especially for language pairs whose word order is different (e.g., English and Japanese), to shorten delay and reduce cognitive load.
-Because of these features, SI sentences are different from offline translation sentences, but most of the studies in automatic SI Oda et al.
+Maintaining word order in a source language is another important strategy, especially for language pairs whose word order differs (e.g., English and Japanese), to shorten delays and reduce cognitive load.
+Because of these features, SI sentences are different from offline translation sentences, although most automatic SI studies Oda et al.
 (2014); Ma et al.
@@
 (2020); Papi et al.
-(2023) have been used offline translation corpora (e.g., MuST-C; Di Gangi et al., 2019) for both training and evaluation of the models due to the limited amount of simultaneous interpretation corpora (SICs).
-For the English-Japanese language pair, several SICs have been constructed Toyama et al.
+(2023) have used offline translation corpora (e.g., MuST-C; Di Gangi et al., 2019) for both training and evaluatng models due to the limited amount of simultaneous interpretation corpora (SICs).
+For English-Japanese language pairs, several SICs have been constructed Toyama et al.
 (2004); Shimizu et al.
@@
 (2021).
-Based on NAIST Simultaneous Interpretation Corpus (NAIST-SIC; Doi et al., 2021), Zhao et al.
-(2024)111The dataset was released in 2023 (see the version 3 of the paper).
-created an automatically-aligned parallel SI dataset, named NAIST-SIC-Aligned.
-Since the sentences in the dataset are aligned at the sentence level, they can be used for model training.
+Based on the NAIST Simultaneous Interpretation Corpus (NAIST-SIC; Doi et al., 2021), Zhao et al.
+(2024)111The dataset was released in 2023 (see version 3 of the paper).
+created an automatically-aligned parallel SI dataset: NAIST-SIC-Aligned.
+Since its sentences are aligned at the sentence level, they can be used for model training.
 Actually, Ko et al.
@@
 (2024) trained SI models using SI data from NAIST-SIC-Aligned.
-The model performance was evaluated through automatic evaluation metrics such as BLEU Papineni et al.
+Their model performances were evaluated through automatic evaluation metrics such as BLEU Papineni et al.
 (2002) using a small test set curated based on SI sentences generated by professional human simultaneous interpreters.
-The scores reported in Ko et al.
+Although the scores reported in Ko et al.
 (2023) and Zhao et al.
-(2024) were relatively low but the test set used in the two studies might have underestimated the model performance.
-Since human simultaneous interpreters use strategies such as summarization and generalization, phrases that do not affect the main idea are not necessarily translated into the target language.
-If an SI model generates translation for the phrases that a human interpreter did not, the output sentence might not be evaluated properly, even when it is correct translation.
+(2024) were relatively low, the test set used in both studies might have underestimated the model performance.
+Since human simultaneous interpreters use such strategies as summarization and generalization, phrases that do not affect the main idea are not necessarily translated into the target language.
+If an SI model generates translations for phrases that a human interpreter did not, the output sentence might not be evaluated properly, even when it is a correct translation.
 | Source | (1) The US Secret Service, / (2) two months ago, / (3) froze the Swiss bank account / (4) of Mr.
@@
 Fukuda et al.
-(2024) pointed out the difficulty for SI models to learn which phrases in source speech are less important and called for constructing SI models that only employ the strategy of maintaining the word order in a source language.
-As the first step, they created NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset222https://dsc-nlp.naist.jp/data/NAIST-SIC/Aligned-Chunk_Mono-EJ/.
+(2024) pointed out the difficulty for SI models to learn which phrases in source speech are less important and advocated constructing SI models that only employ a strategy that maintains the word order in a source language.
+As a first step, they created the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset222https://dsc-nlp.naist.jp/data/NAIST-SIC/Aligned-Chunk_Mono-EJ/.
 The source sentences in the test set used in Ko et al.
-(2023) were automatically segmented into chunks, and each chunk was translated in a way that did not include the content of the subsequent chunks.
-Unlike in SI sentences by human interpreters, where not all information in the source sentences is translated, chunk-wise monotonic translation (CMT) was developed so that all information is translated (Table 1)333Precisely, omission to maintain the fluency of the sentence was allowed.
+(2023) were automatically segmented into chunks, each of which was translated in a way that did not include the content of subsequent chunks.
+Unlike in SI sentences by human interpreters, where not all the information in the source sentences is translated, chunk-wise monotonic translation (CMT) sentences333CMT refers to the task of segmenting a source sentence into chunks and translating it in the order of the chunks.
+A CMT sentence is a target sentence generated through CMT.
+were translated so that all the information is translated (Table 1)444Precisely, omissions that maintained the fluency of the sentence were allowed.
 See Section 3.1 for the details about the dataset..
 Fukuda et al.
-(2024) have investigated the quality of CMT through human evaluation, but have not analyzed the characteristics of it.
-In addition, they have not conducted any evaluation experiments in which model outputs are evaluated on the CMT Evaluation dataset.
-In this paper, we analyze CMT sentences qualitatively and quantitatively.
-In the process of developing CMT, it was allowed to repeat, defer, and omit phrases in the source sentences to keep the translation fluent.
-We suppose that there are factors (e.g., syntactic structures) that prevent monotonic translation if phrases were repeated, deferred, or omitted in CMT sentences since the sentences were translated without time constraints.
-In addition, we evaluate the output from (1) a speech translation (ST) model trained on offline data, (2) a simultaneous speech translation (simulST) model trained on offline data, and (3) a simulST model trained on both offline and SI data.
+(2024) have investigated the quality of the CMT sentences in their dataset through human evaluation, although they have not analyzed its characteristics.
+Nor have they conducted any evaluation experiments in which model outputs are evaluated on their dataset.
+In this paper, we qualitatively and quantitatively analyze CMT sentences in the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset.
+In the process of generating CMT sentences for the dataset, it was allowed to repeat, defer, and omit phrases in the source sentences to maintain the translation’s fluency.
+We assume the presence of factors (e.g., syntactic structures) that prevent monotonic translation if phrases were repeated, deferred, or omitted in the CMT sentences since they were translated without time constraints.
+In addition, we evaluate the output from an existing speech translation (ST) model and two simultaneous speech translation (simulST) models (See 5.2).
+Both the ST555A ST model generates translations after the utterances are completed.
+and simulST models are used to investigate the differences in scores when evaluating translations with different characteristics.
 The contributions of this paper are as follows:
-We analyze CMT sentences and show that they tend to be longer than offline translation primarily because of repetition.
-We investigate what makes the phrases in source sentences to be repeated, deferred, and omitted, and show that most cases occur because of particular grammatical structures.
+We analyze CMT sentences and show that they tend to be longer than offline translations primarily because of repetition.
+We investigate what causes the phrases in source sentences to be repeated, deferred, and omitted and show that most cases occur because of particular grammatical structures.
 When a phrase in a chunk is a dependent of a phrase in the preceding chunk, the head phrase is typically repeated or deferred.
-We evaluate the output from three different models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset: (1) an ST model trained on offline data, (2) a simulST model trained on offline data, and (3) a simulST model trained on both offline and SI data.
-We show that the existing SI-based test set Ko et al.
+We evaluate the output from three different models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset: (1) an ST model trained on offline data, (2) a simulST model trained on offline data, and (3) a simulST model trained on both offline and SI data.
+The results suggest that the existing SI-based test set Ko et al.
 (2023); Zhao et al.
 (2024) underestimates the model performance.
-The results also show that the CMT evaluation dataset would better evaluate simulST models while using an offline-based test set for evaluating simulST models underestimates the model performance.
+The results also suggest that using CMT sentences as references gives higher scores to simulST models than ST models, while using an offline-based test set for evaluating simulST models underestimates the model performance.
 2 Related Work
 2.1 Simultaneous Interpretation Corpora
-SICs are valuable resources both for developing automatic SI models and for analyzing the characteristics of SI.
-For the English-Japanese language pair, several SICs are publicly available Toyama et al.
+SICs are valuable resources both for developing automatic SI models and analyzing SI’s characteristics.
+For English-Japanese language pairs, several SICs are publicly available Toyama et al.
 (2004); Shimizu et al.
@@
 (2020); Doi et al.
-(2021) although the amount of such corpora is very limited compared to offline translation corpora.
+(2021), although the amount of such corpora is very limited compared to offline translation corpora.
 Using these corpora, SI sentences have been analyzed from various perspectives, such as strategies and interpreting patterns used by interpreters, latency, translation quality, and word order Tohyama and Matsubara (2006); Ono et al.
@@
 (2021).
-Furthermore, SI models have also been developed using SICs Ryu et al.
+SI models have also been developed using SICs Ryu et al.
 (2004); Shimizu et al.
@@
 2.2 Word Order in Simultaneous Interpreting
-When dealing with language pairs whose sentence structures are different, including English and Japanese (SVO and head-initial vs.
-SOV and head-final), reducing the word order difference between the source and the target language is important to minimize the delay.
+When dealing with language pairs whose sentence structures are different, including English and Japanese (SVO/head-initial vs.
+SOV/head-final), reducing the word order differences between the source and the target languages is crucial for minimizing delays.
 Murata et al.
-(2010) segmented source sentences into semantically meaningful units with a maximum length of 4.3 seconds, and translated those units from the viewpoint of SI.
+(2010) segmented source sentences into semantically meaningful units with a maximum length of 4.3 seconds and translated those units from an SI viewpoint.
 He et al.
@@
 (2020) reordered Japanese sentences to make the word order closer to the original English sentences.
-They further applied style transfer to make the sentence fluent and obtained sentences close to SI sentences by human interpreters.
+They further applied style transfer to increase the fluency and obtained sentences close to SI sentences by human interpreters.
 Han et al.
@@
 They trained SI models for four language pairs, including English-Japanese.
-Nakabayashi and Kato (2021) segmented sentences into chunks and created bilingual pairs of the chunks with explicit annotations of context information.
-The SI model trained on the data translated the source sentences while referencing preceding chunks but connecting chunks naturally remained a challenge.
+Nakabayashi and Kato (2021) segmented sentences into chunks and created bilingual pairs of such chunks with explicit annotations of context information.
+The SI model trained on the data translated the source sentences while referencing the preceding chunks although naturally connecting chunks remained a challenge.
 Higashiyama et al.
-(2023) have constructed a large-scale English Japanese SIC with information of chunk boundaries in source and target sentences and phrases that can be omitted in target sentences.
-NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset Fukuda et al.
-(2024), which is similar to Higashiyama et al.’s (2023), is relatively small and intended for the evaluation purpose.
+(2023) constructed a large-scale English Japanese SIC with the information of chunk boundaries in source and target sentences and phrases that can be omitted in target sentences.
+The NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset Fukuda et al.
+(2024), which is similar to Higashiyama et al.’s (2023), is relatively small and intended for the evaluation purposes.
 The word order differences among different translation modes have also been investigated.
@@
 While Cai et al.
-(2020) analyzed actual SI data generated by human simultaneous interpreters, we use the CMT dataset, which is created without time constraints, and in which all information in the source sentences is translated into the target sentences.
+(2020) analyzed the actual SI data generated by human simultaneous interpreters, we use CMT sentences, which were generated without time constraints, and in which all the information in the source sentences is translated into target sentences.
 3 Chunk-wise Monotonic Translation
-In SI between language pairs with different sentence structures, interpreters are known to segment source sentences into chunks and translate them in the order of the chunks444Reordering may occur within a chunk.
+In SI between language pairs with different sentence structures, interpreters segment source sentences into chunks and translate them from chunk to chunk666Reordering may occur within a chunk.
 Okamura and Yamada (2023).
-This section describes the details of NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset, which is used in our analyses.
+This section describes the details of the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset, which is used in our analyses.
 3.1 Data
-NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset consists of 511 pairs of source sentences and their corresponding chunk-wise monotonic translation (CMT) sentences with information of chunk boundaries in source and target sentences.
+The NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset consists of 511 pairs of source sentences and their corresponding chunk-wise monotonic translation (CMT) sentences with information of chunk boundaries in the source and target sentences.
 Source (i.e., English) sentences, which were used as the test set in Ko et al.
-(2023), were segmented following the five rules that reflected interpreter’s strategies555See the original paper or the code for chunking: https://github.com/ahclab/si_chunker based on the syntactic analysis results from spaCy.
-The source sentences come from 8 TED talks.
-| Data | Sum | Per Sent. |
+(2023), were segmented following the five rules that reflected the interpreter’s strategies777See the original paper or the code for chunking: https://github.com/ahclab/si_chunker based on the syntactic analysis results from spaCy.
+The source sentences come from eight TED talks.
+| Data | Sum | Per Sent.SD |
 | # sentence pairs | 511 | – |
-| # chunks | 1,677 | 3.28 |
-| # source words | 8,104 | 15.86 |
-| # target words | 13,981 | 27.36 |
-Translators were provided with source sentences with chunk boundaries and were asked to (1) translate them in the order of the chunks while (2) connecting the chunks naturally and (3) not including the content of the subsequent chunks.
-It was allowed to (1) repeat, (2) defer, and (3) omit phrases in the source sentences to keep the translation fluent, but was instructed to avoid using the operations with larger number as much as possible.
+| # chunks | 1,677 | 3.282.12 |
+| # source words | 8,104 | 15.8610.16 |
+| # target words | 13,981 | 27.3618.55 |
+Translators were provided with source sentences with chunk boundaries and asked to (1) translate them in the order of the chunks while (2) naturally connecting the chunks and (3) not including the content of subsequent chunks.
+They were allowed to (1) repeat, (2) defer, and (3) omit phrases in the source sentences to keep translation fluency, although they were instructed to minimize their use of the operations with larger number as much as possible (e.g., defer should not be used when repeat can handle the situation) .
 Data statistics are shown in Table 2.
@@
 (2024) have examined the quality of CMT sentences through human evaluation but have not analyzed the characteristics of them.
-We suppose that there are factors that prevent monotonic translation if a phrase in the source sentences is repeated, deferred, and omitted since the CMT sentences were generated without time constraints.
-We qualitatively and quantitatively analyze the CMT sentences with these operations and reveal the factors.
-To better understand the characteristics of CMT sentences, we also compare them with SI sentences from NAIST-SIC and NAIST-SIC-Aligned, and offline translation sentences from the subtitles of TED talks.
-Since the SI sentences in NAIST-SIC were not aligned at the sentence level, we performed alignment manually.
-Some source sentences did not match across the datasets, and we excluded those 10 sentences from the analyses.
-In addition, 25 sentences were not translated in NAIST-SIC666For example, due to time constraints, interpreters may not be able to translate a whole sentence.
+We suppose that factors exist that prevent monotonic translation if a phrase in the source sentences is repeated, deferred, or omitted since the CMT sentences were generated without time constraints.
+Therefore we qualitatively and quantitatively analyze the CMT sentences with these operations and reveal such factors.
+To better understand the characteristics of the CMT sentences, we also compare them with SI sentences from NAIST-SIC and NAIST-SIC-Aligned as well as offline translation sentences from the subtitles of TED talks.
+Since the SI sentences in NAIST-SIC were not aligned at the sentence level, we manually align them.
+Some source sentences did not match across the datasets, and we excluded those ten sentences from the analyses.
+In addition, 25 sentences were not translated in NAIST-SIC888For example, due to time constraints, interpreters might have been unable to translate a whole sentence.
 See Doi et al.
@@
 4.1 Annotations
-To analyze the characteristics of CMT sentences, we annotated tags to source and CMT sentences.
+To analyze the characteristics of the CMT sentences, we annotated tags to the source and CMT sentences.
 The list of tags is shown in Table 3.
-Prior to the annotation, we tokenized the English sentences using spaCy777https://spacy.io/ and the Japanese sentences using MeCab Kudo et al.
+Prior to the annotations, we tokenized the English sentences using spaCy999https://spacy.io/ and the Japanese sentences using MeCab Kudo et al.
 (2004) with unidic.
-Then, the source and CMT sentences were concatenated with special token [SEP] and annotated using an open-source data labeling tool, doccano888https://github.com/doccano/doccano.
+Then, we concatenated the source and CMT sentences with a special token [SEP] and annotated them using an open-source data labeling tool, doccano.101010https://github.com/doccano/doccano
+We identified spans (i.e., words or phrases) that are repeated, deferred, or omitted and annotated the span tags.
+In addition, we annotated ahead, add, and error tags for analyses of problematic translations.
+The corresponding span tags in the source and CMT sentences were associated using relation tags.
 Annotation examples are shown in Figure 1.
-The first and second authors collaboratively annotated the first 50 examples while discussing. As sufficient agreement was expected, the remainder of the data were annotated by the first author.
-4.2 Results of Analyses
+The first and second authors collaboratively annotated the first 50 examples while discussing their decisions.
+Since sufficient agreement was assumed, the remainder of the data were just annotated by the first author.
+4.2 Analysis Results
 4.2.1 Comparison among Different Translation Modes
-We compared the length of sentences in the four datasets.
+We compared the sentence lengths of the four datasets.
 Fukuda et al.
-(2024) have also conducted similar comparisons based on the number of characters.
-However, there were variations in spelling and differences in transcription systems (e.g., numbers), and we made comparisons based on the number of words segmented by MeCab.
-Table 4 show that CMT sentences were the longest, followed by offline, NAIST-SIC, and NAIST-SIC-Aligned, which matched the results reported in Fukuda et al.
+(2024) also conducted similar comparisons based on the number of characters.
+However, since variations in spelling and differences in transcription systems (e.g., numbers) were found, we made comparisons based on the number of words segmented by MeCab.
+Table 4 shows that the CMT sentences were the longest, followed by offline, NAIST-SIC, and NAIST-SIC-Aligned.
+These results matched those reported in Fukuda et al.
 (2024).
-Long translation sentences can pose some problems.
+Long translated sentences can pose some problems.
 As discussed in Fukuda et al.
-(2024), it may increase the cognitive load on the listeners/readers of the translations.
+(2024), the lenght may increase the cognitive load on the listeners/readers of the translations.
 In addition, longer output may cause a delay even though CMT aims to reduce it.
-| Dataset | Sum | Per Sent. |
-| CMT | 13,508 | 28.38 |
-| NAIST-SIC | 8,914 | 18.73 |
-| NAIST-SIC-Aligned | 8,072 | 16.96 |
-| Offline | 9,907 | 20.81 |
-4.2.2 Factors that Make CMT Sentences Longer
-To reveal what made the CMT sentences longer, we first analyzed them qualitatively.
-Our analyses suggest that (1) CMT sentences contain many formulaic expressions for the end of sentences as sentences are segmented into small chunks, and (2) words that are often omitted in Japanese (e.g., pronouns) are explicitly translated since translators were instructed to avoid omitting phrases in the source language, as in the following examples.
+| Dataset | Sum | Per Sent.SD |
+| CMT | 13,508 | 28.3818.66 |
+| NAIST-SIC | 8,914 | 18.7312.08 |
+| NAIST-SIC-Aligned | 8,072 | 16.9611.52 |
+| Offline | 9,907 | 20.8112.62 |
+4.2.2 Factors that Lengthen CMT Sentences
+To reveal what factor lengthened the CMT sentences, we first analyzed them qualitatively.
+Our analyses suggest that (1) CMT sentences contain many formulaic expressions for the end of sentences as they are segmented into small chunks, and (2) the words that are often omitted in Japanese (e.g., pronouns) are explicitly translated since translators were instructed to avoid omitting phrases in the source language, as in the following examples:
 (En) It’s when we warmed it up, / and we turned on the lights / and looked inside the box, / we saw that the piece metal / was still there in one piece.
@@
 (offline) 目だけしか動かせません
-In addition to the above characteristics, there were many repetitions observed in particularly long sentences.
-To verify this, we further analyzed particularly long and short sentences, chosen based on the length ratio of CMT sentences to offline ones.
+In addition to the above characteristics, we observed many repetitions in particularly long sentences.
+To verify this, we further analyzed particularly long and short sentences, chosen based on the length ratio of the CMT sentences to the offline ones.
 The long and short sentences were defined as those with a length ratio greater/smaller than the average 0.5 standard deviations (avg.=1.39, SD=0.43).
-We subjectively judged if those sentences contained many repetitions.
-We also identified the sentences whose offline translations were short.
+We subjectively judged whether these sentences contained many repetitions.
+We also identified sentences whose offline translations were short.
 Table 5 shows that long sentences contained more repetitions than short ones.
 The offline translation sentences were short, probably because they were originally subtitles, for which limited space was allowed.
-We also checked it quantitatively using the number of repeat tags assigned to the sentences, and found that the frequency of repetition tags was higher in long sentences (Table 5).
+We also quantitatively checked them using the number of assigned repeat tags and found that the frequency of repetition tags was higher in long sentences (Table 5).
 | Type | N | Repeat | Repeat | Short |
@@
 4.2.3 Omission in SI Sentences
-To find techniques to shorten translations, we analyzed SI sentences in NAIST-SIC.
-Based on the length ratio of SI sentences to offline ones, we defined SI sentences that potentially involved reasonable omissions (omission; the ratio ), and SI sentences that may not fully convey the meaning of the source sentences (undertranslation; the ratio ), following the criteria in Higashiyama et al.
+To find techniques for shortening translations, we analyzed the SI sentences in NAIST-SIC.
+Based on the length ratio of SI sentences to the offline ones, we defined SI sentences that might have reasonable omissions (omission; ratio ) and SI sentences that probably failed to fully convey the meaning of the source sentences (undertranslation; ratio ), following the criteria in Higashiyama et al.
 (2023).
-We expected to find some trends (e.g., part-of-speech) in phrases that were omitted, but we was not able to.
+Although we expected to identify some trends (e.g., part-of-speech) in the phrases that were omitted, we did not do so.
 In addition, we found a certain number of unacceptable translations in both categories (43.12% and 60.00% for omission and undertranslation, respectively).
-The results suggest that human simultaneous interpreters judge the importance of the phrases based on context and decide whether to translate them, which are sometimes correct and sometimes not.
+The results suggest that human simultaneous interpreters judge the importance of phrases based on context and decide whether to translate them; some judgements are correct, and some are not.
 4.2.4 Factors that Make Monotonic Translation Difficult
-With the help of tags annotated to source and CMT sentences, we analyzed factors that make monotonic translation difficult.
-Table 6 shows the number of source phrases that were repeated, deferred, and omitted.
-The values are based on the number of rel-repeat, defer_d#, and omit tags, respectively.
-We counted the number of relation tags for repeat and defer because the span tags for the two operations were assigned to both source and CMT sentences.
+With the help of tags annotated to the source and CMT sentences, we analyzed the factors that make monotonic translation difficult.
+Table 6 shows the number of source phrases that were repeated, deferred, or omitted.
+The values are based on the number of rel-repeat, defer_d#, and omit tags.
+We counted the relation tags for repeat and defer because the span tags for those two operations were assigned to both the source and CMT sentences.
 The results show that the translators used repeat most frequently, followed by defer and omit, as they were instructed (see Section 3.1).
@@
 For these phrases, we explored what makes monotonic translations difficult.
-Our analyses revealed that most cases of repeat and defer occurred because of particular grammatical structures.
-Table 7 lists major structures along with their frequencies in the data and examples.
+Our analyses revealed that most cases of repeat and defer were caused by particular grammatical structures.
+Table 7 lists the major structures along with their frequencies in the data and examples.
 In these structures, a phrase in a chunk is typically a dependent of a phrase in the preceding chunk.
 In the example of a post-modifier (Table 7), the relative pronoun clause is a dependent of the noun phrase a device, which is in a preceding chunk.
-When phrases with a dependency relation exist across multiple chunks, CMT would be difficult because Japanese is a strongly head-final language.
-The examples in Table 7 show how human translators address these structures by repeating or deferring some phrases in the subsequent chunks.
+When phrases with a dependency relation exist across multiple chunks, CMT is difficult because Japanese is a strongly head-final language.
+The examples in Table 7 show how human translators address these structures by repeating or deferring some phrases in subsequent chunks.
 Prepositions, post-modifiers, and dependent clauses have also been identified as syntactic factors that affect interpreters’ word order decisions in Cai et al.
 (2020).
-Human interpreters find those structures challenging for SI and adopt a strategy to maintain the word order of the source language.
+Human interpreters find these structures challenging for SI and adopt a strategy to maintain the word order of the source language.
 | Structures | # repeat | # defer | Examples |
@@
 | {CJK} UTF8min … / 全体主義政府にツールを提供しているということです、 / これを行うための、 / 自国の市民に対してこれを行うためのツールを。 [providing totalitarian governments with tools / to do this / tools to do this against their own citizens] |
-In addition, we observed that inappropriate segmentation was addressed through repeating and deferring the phrases.
-Most of the inappropriate segmentation was found in phrasal verbs, -ing forms of verbs, and to-infinitives.
-In SI data, we also found human interpreters repeat phrases to maintain the word order of the source language.
-For example, in the example in Table 1, a noun modified by a preposition phrase is repeated.
+In addition, we observed that inappropriate segmentation was addressed by repeating and deferring the phrases.
+Most inappropriate segmentation was found in phrasal verbs, verbal gerunds, and to-infinitives.
+In the SI data, we also found that human interpreters repeat phrases to maintain the word order of the source language.
+For example, in the example in Table 1, a noun modified by a preposition phrase is repeated:
 (En) … / froze the Swiss bank account / of Mr. Sam Jain right here, / …
@@
 [… / froze the Swiss bank account / it is Sam Jain’s one / …]
-In addition, Okamura and Yamada (2023) reported that the order of the chunks is swapped about once on average in an SI sentence.
-These things suggest that human interpreters address the word order differences that make monotonic translation difficult through repeating and deferring some phrases.
-5 Evaluation Using CMT Data
-To investigate the impact of using CMT sentences for evaluating the translation quality, we evaluated the output from existing ST and simulST models using NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as existing test sets.
+In addition, Okamura and Yamada (2023) reported that the order of the chunks is shuffled about once on average in an SI sentence.
+These things suggest that human interpreters address the word order differences that make monotonic translation difficult by repeating and deferring some phrases.
+5 Evaluation Using CMT sentences
+To investigate the impact of using CMT sentences for evaluating translation quality, we evaluated the output from existing ST and simulST models using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as existing test sets.
 5.1 Data
-We used the following datasets as the reference for the automatic evaluation metrics:
-n-cmt: CMT sentences from NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset
-si_hum: SI sentences from NAIST-SIC, which were manually aligned to the source speech
-si_auto: SI sentences from NAIST-SIC-Aligned, which were aligned automatically
+We used the following four datasets as references for the automatic evaluation metrics:
+n-cmt: CMT sentences from the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset
+si_hum: SI sentences from NAIST-SIC, manually aligned to the source speech
+si_auto: SI sentences from NAIST-SIC-Aligned, aligned automatically
 offline: offline translation sentences from the subtitles of TED talks.
+Because si_auto was created by applying automatic alignment and filtering techniques to SI sentences in si_hum, it may contain alignment errors.
+In addition, SI sentences in si_auto tend to be shorter than those in si_hum (see Table 4).
+Therefore, we used the two SI-based datasets for our evaluation.
 5.2 Speech Translation Models
-We used three existing ST and simulST models:
+We used three existing models (i.e., one ST and two simulST models):
 ST_offline: an ST model trained on offline data Fukuda et al. (2023)
 simulST_offline: a simulST model trained on offline data Ko et al. (2023)
-simulST_si_offline: a simulST model trained on both offline and SI data Ko et al. (2023)
+simulST_si_offline: a simulST model trained on both offline and SI data Ko et al. (2023).
 All the models were built by connecting two pre-trained models, HuBERT-Large Hsu et al.
@@
 (2020) for their text decoder.
-The encoder and decoder were connected via Inter-connection Nishikawa and Nakamura (2023) and a length adapter Tsiamas et al.
+The encoder and decoder were connected by Inter-connection Nishikawa and Nakamura (2023) and a length adapter Tsiamas et al.
 (2022).
@@
 (2020).
-For ST_offline, we used the model with checkpoint averaging (Inter-connection + Ckpt Ave.
+For ST_offline, we used a model with checkpoint averaging (Inter-connection + Ckpt Ave.
 in Fukuda et al.
 (2023)).
-For simulST_offline and simulST_si_offline, we used the models which satisfy the task requirement of the simultaneous track in IWSLT 2023 Evaluation Campaign999https://iwslt.org/2023/simultaneous, latency measured by Average Lagging Ma et al.
+For simulST_offline and simulST_si_offline, we used the models that satisfy the task requirement of the simultaneous track in the IWSLT 2023 Evaluation Campaign111111https://iwslt.org/2023/simultaneous, latency measured by Average Lagging Ma et al.
 (2019) seconds (Offline FT and Mixed FT + Style in Ko et al.
 (2023), respectively).
-5.3 Metrics
-We evaluated the translation quality of the output from the ST and simulST models (see Section 5.2) using BLEU101010BLEU was calculated using sacreBLEU.
-Post (2018) https://github.com/mjpost/sacrebleu, BLEURT111111https://github.com/google-research/bleurt Sellam et al.
-(2020), COMET121212https://github.com/Unbabel/COMET Rei et al.
-(2020), and BERTScore131313https://github.com/Tiiiger/bert_score Zhang et al.
-(2020).
-BERTScore was calculated using bert-base-multilingual-cased.
-We used the four datasets described in Section 5.1 as the reference.
-5.4 Evaluation Results
 | Model | BLEU | BLEURT | COMET | |||||||||
 | n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | n-cmt | si_hum | si_auto | offline | |
-| ST_offline | 14.487 | 8.856 | 8.637 | 17.775 | 0.553 | 0.447 | 0.414 | 0.538 | 0.838 | 0.797 | 0.781 | 0.833 |
-| simulST_offline | 15.406 | 8.446 | 7.773 | 17.907 | 0.556 | 0.442 | 0.406 | 0.531 | 0.826 | 0.780 | 0.763 | 0.821 |
-| simulST_si_offline | 15.982 | 12.031 | 11.020 | 13.191 | 0.567 | 0.493 | 0.460 | 0.519 | 0.807 | 0.774 | 0.761 | 0.789 |
+| ST_offline | 14.487 | 8.856 | 8.637 | 17.775 | 0.553 | 0.447 | 0.414 | 0.538 | 0.838 | 0.797 | 0.781∗1 | 0.833 |
+| simulST_offline | 15.406† | 8.446† | 7.773† | 17.907 | 0.556 | 0.442 | 0.406 | 0.531 | 0.826 | 0.780 | 0.763 | 0.821 |
+| simulST_si_offline | 15.982† | 12.031† | 11.020† | 13.191† | 0.567 | 0.493∗1 | 0.460∗1 | 0.519 | 0.807∗2 | 0.774∗3 | 0.761 | 0.789∗2 |
 | Model | BERTScore (Pre.) | BERTScore (Rec.) | BERTScore (F1) | |||||||||
@@
 | simulST_offline | 0.799 | 0.730 | 0.717 | 0.783 | 0.770 | 0.738 | 0.734 | 0.786 | 0.783 | 0.734 | 0.725 | 0.784 |
-| simulST_si_offline | 0.817 | 0.764 | 0.746 | 0.759 | 0.784 | 0.766 | 0.760 | 0.757 | 0.800 | 0.764 | 0.752 | 0.757 |
-Table 8 shows the results of quality evaluation metrics across ST and simulST models.
+| simulST_si_offline | 0.817∗1 | 0.764∗1 | 0.746∗1 | 0.759∗2 | 0.784∗1 | 0.766∗1 | 0.760∗1 | 0.757∗2 | 0.800∗1 | 0.764∗1 | 0.752∗1 | 0.757∗2 |
+5.3 Metrics
+We evaluated the translation quality of the output from the ST and simulST models (see Section 5.2) using BLEU121212BLEU was calculated using sacreBLEU.
+Post (2018) https://github.com/mjpost/sacrebleu, BLEURT131313https://github.com/google-research/bleurt Sellam et al.
+(2020), COMET141414https://github.com/Unbabel/COMET Rei et al.
+(2020), and BERTScore151515https://github.com/Tiiiger/bert_score Zhang et al.
+(2020).
+BERTScore was calculated using bert-base-multilingual-cased.
+We used the four datasets described in Section 5.1 as references.
+5.4 Evaluation Results
+Table 8 shows the results of the quality evaluation metrics across the ST and simulST models.
+For the BLEU scores, we conducted paired significance tests using paired bootstrap resampling Koehn (2004).
+We specified ST_offline as the baseline for the significance tests.
+For the other scores, we conducted a one-way ANOVA, followed by Tukey’s multiple comparisons test.
 When the translation quality was evaluated using BLEU with n-cmt as the reference, simulST_si_offline achieved the highest score.
 On the SI-based test sets (i.e., si_hum and si_auto), simulST_si_offline also had the highest score.
-On the offline-based test set, in contrast, models trained on only offline data achieved much higher score than simulST_si_offline.
+On the offline-based test set, in contrast, the models trained on only offline data achieved much higher scores than simulST_si_offline.
 The same tendencies were observed in BLEURT and BERTScore.
-These results suggest that the model trained on both SI and offline data generated more SI-like translations, and that such models should be evaluated using a reference closer to SI sentences.
-In addition, using an offline-based test set might underestimate the performance of the model trained on both SI and offline data.
-Comparing n-cmt, si_hum, and si_auto, the scores were highest for n-cmt, followed by si_hum and si_auto on all metrics and models.
-Because si_hum is based on SI sentences generated by human simultaneous interpreters, some content in a source speech might be omitted or inadequately translated (under-translation).
+These results suggest that the models trained on both SI and offline data generated more SI-like translations, and such models perhaps should be evaluated using a reference closer to SI sentences.
+In addition, using an offline-based test set might underestimate the performance of models trained on both SI and offline data.
+Comparing n-cmt, si_hum, and si_auto, the scores were highest for n-cmt, followed by si_hum and si_auto on all the metrics and models.
+Because si_hum is based on SI sentences generated by human simultaneous interpreters, some content in the source speech might be omitted or inadequately translated (under-translation).
 SI sentences in si_auto, which were automatically created based on human SI sentences, might contain less source speech content than those in si_hum due to the alignment and filtering techniques applied (see Zhao et al., 2024).
-In fact, BERTScore precision was higher than recall on n-cmt, in which there are almost no omissions, while the recall was higher than the precision on si_auto and the precision and the recall were almost equal on si_hum.
-These results suggest that the existing SI-based test sets Ko et al.
+In fact, BERTScore precision was higher than recall on n-cmt, in which there were almost no omissions, while recall was higher than precision on si_auto and precision and recall were almost equal on si_hum.
+These results indicate the possibility that the existing SI-based test sets Ko et al.
 (2023); Zhao et al.
 (2024) underestimate the model performance.
-However, the results on COMET were different from those on other metrics (Table 8).
+However, the COMET results were different from those on the other metrics (Table 8).
 On all four test sets, ST_offline achieved the highest score, followed by simulST_offline and simulST_si_offline.
-One of the possible reasons is that COMET uses source sentences to calculate scores.
-To examine the impact of source sentences, we also calculated the reference-free COMET-QE using wmt22-cometkiwi-da and got similar results (0.813, 0.798, 0.766 for ST_offline, simulST_offline, simulST_si_offline, respectively).
+One possible reason is that COMET uses source sentences to calculate its scores.
+To examine the impact of the source sentences, we also calculated a reference-free COMET-QE using wmt22-cometkiwi-da and got similar results (0.813, 0.798, and 0.766 for ST_offline, simulST_offline, and simulST_si_offline, respectively).
 We further calculated COMET-QE for n-cmt and offline, regarding them as oracle data, and found that n-cmt had a higher score than offline (n-cmt: 0.832, offline: 0.812).
-Because some translation sentences in offline include under-translation, these results suggest that the COMET score tends to become higher when more content in the source sentences is covered in the target sentences.
+Because some translation sentences in offline are under-translated, these results suggest that the COMET scores tend to become high when more content in the source sentences is covered in the target sentences.
 This feature does not fit the nature of SI, where human interpreters use sophisticated strategies (see He et al., 2016; Cai et al., 2020, for example).
-We need to carefully interpret the COMET score when we use it for evaluating simulST models.
+We need to carefully interpret COMET scores when we use them for evaluating simulST models.
 6 Conclusion
-This paper focuses on monotonic translations in English-Japanese SI.
-Our analyses revealed some syntactic factors that make monotonic translations difficult and that human interpreters/translators address these challenges by repeating or deferring some phrases in the source language in the subsequent chunks.
-The grammatical structures that might cause delay would be useful information for developing segmentation or decoding policies of simultaneous machine translation systems.
-We also evaluated the output from the existing ST and simulST models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing SI-based and offline-based test sets.
-The results suggest that monotonic translation data would better evaluate simulST models trained using SI data, and that the existing test sets underestimate the performance of such models.
-A future direction would be using monotonic translation data for developing simulST models.
+This paper focused on monotonic translations in English-Japanese SI.
+Our analyses revealed some grammatical structures that make monotonic translations difficult and that human interpreters/translators address these challenges by repeating or deferring some phrases in source language in the subsequent chunks.
+The grammatical structures that might cause delays would be useful information for developing segmentation or decoding policies for simultaneous machine translation systems.
+One possible direction would be predicting whether a phrase in a chunk is the head of a phrase in subsequent chunks.
+We also evaluated the output from the existing ST and simulST models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing SI-based and offline-based test sets.
+The BLEU, BLEURT, and BERTScore results supported using CMT sentences for evaluating simulST models trained using SI data, although the results with COMET were different.
+Further analysis across various evaluation metrics is necessary.
+Analyzing how the source and target sentences are aligned monotonically on different types of translations (e.g., Han et al., 2021) would also be useful.
+This paper investigated the impact of using CMT sentences for evaluation purposes.
+A future study would involve using monotonic translation sentences for developing simulST models Sakai et al.
+(2024)161616Published around the same time as the submission of this paper..
+It could potentially address the problem that simulST models trained using SI sentences suffered from under-translation Ko et al.
+(2023).
+However, CMT sentences tend to be long.
+Investigating the trade-offs between longer CMT sentences and the potential cognitive load on listeners/readers might provide further insights.
 Acknowledgments